OpenStudy (anonymous):

4x – (6 – x) = 12 + 4x A. –18 B. 2 C. –2 D. 18

OpenStudy (anonymous):

plz help me

OpenStudy (anonymous):

ill give a medal

OpenStudy (anonymous):

im working it out one sec

OpenStudy (anonymous):

ok

jagr2713 (jagr2713):

Its D

OpenStudy (anonymous):

989 i got another problem for you

OpenStudy (anonymous):

um dude ur not suppose to just give out the anwser u have to explain

OpenStudy (anonymous):

ill get you a medal too

OpenStudy (anonymous):

9(x – 2) = 5(2 + x)

OpenStudy (anonymous):

i need this one solved

jagr2713 (jagr2713):

u need to distribute

jagr2713 (jagr2713):

9x-18=10+5x

OpenStudy (neer2890):

9x-18=10+5x 9x-5x=18+10 4x=28 x=7

jagr2713 (jagr2713):

wowo yea 7

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

9x-18=10+5x you need to put x's on one side and numbers on the other subtract 9 by 5 4x-18=10 add 18 to 10 4x=28 divide 28 by 4 x=7